# Long time asymptotics for the KPII equation

### v3

@@ 3.1. Estimates on scattering data
(3.2) $\begin{gathered}G_{u_{0}}f=1+\iint e^{2\pi i(x_{1}\xi_{1}+x_{2}\xi_{2})}\frac{\left[u_{0}f\right]^{\wedge}(\xi_{1},\xi_{2},\lambda)}{p_{\lambda}(\xi_{1},\xi_{2})}d\xi_{1}d\xi_{2},\\ p_{\lambda}(\xi_{1},\xi_{2})=(2\pi i\xi_{1}+\lambda)^{2}-(2\pi i\xi_{2}+{\lambda}^{2}).\end{gathered}$
Lemma 3.1.
(3.3) $|G_{u_{0}}f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}|f|_{L^{\infty}},$
(3.4) $|\left[\partial_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}},\ |\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}},$
and, for $j\geq 2$,
(3.5) $|\left[\partial^{j}_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|\lambda_{I}|)^{j-1}(1+|x_{2}|)^{j-1}|u_{0}|_{\mathfrak{M}^{j-1,0}}|f|_{L^{\infty}}+C|\partial^{j-1}_{\lambda_{R}}\left(\lambda_{I}G_{u_{0}}f\right)|,$
(3.6) $|\left[\partial^{j}_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|\lambda|)^{j-1}(1+|x_{2}|)^{j-1}|u_{0}|_{\mathfrak{M}^{j-1,0}}|f|_{L^{\infty}}+C|\partial^{j-1}_{\lambda_{I}}\left(\lambda_{I}G_{u_{0}}f\right)|,$
(3.7) $|\left[\partial_{\lambda_{R}}\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq$ $C(1+|{\color[rgb]{0,0,0}\lambda_{R}}|)(1+|x_{2}|)|u_{0}|_{\mathfrak{M}^{1,0}}|f|_{L^{\infty}}+C|\partial_{\lambda_{I}}\left(\lambda_{I}G_{u_{0}}f\right)|.$
Proof.
We can prove (3.3) by applying the estimates [8]:
@@
where $\Omega_{\lambda}=\{(\xi_{1},\xi_{2})\in{\mathbb{R}}^{2}:|p_{\lambda}(\xi_{1},\xi_{2})|<1\}$.
For $(x_{2}-x_{2}^{\prime})\neq 0$, via the Fourier inversion formula, the convolution, and the residue theorem, the operator $G$ can be written as [1]:
@@
$=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}\left(4\pi i(x_{2}-x_{2}^{\prime})\xi_{1}\right)$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
@@
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}{\color[rgb]{0,0,0}+}\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi\lambda_{I})G_{u_{0}}f.$
Along with (3.3), yields
$|\left[\partial_{\lambda_{R}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}}.$
Similarly,
(3.11) $\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f=$ $-\frac{1}{2\pi i}\iint dx^{\prime}_{1}dx^{\prime}_{2}\ \left[u_{0}f\right](x^{\prime}_{1},x^{\prime}_{2})\int d\xi_{1}e^{2\pi i[(x_{1}-x_{1}^{\prime})+(x_{2}-x_{2}^{\prime})2\lambda_{R}]\xi_{1}}$
@@
$|\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|)|f|_{L^{\infty}}.$
Hence the proof of (3.4) is completed.
Moreover, from (3.10),
$\left[\partial_{\lambda_{R}}^{2}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)^{2}(4\pi\lambda_{I})\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
${\color[rgb]{0,0,0}+}$ $\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi i\lambda_{I})\partial_{\lambda_{R}}\left(Gf\right).$
Hence the proof of (3.5)-(3.7) can be justified by bootstrapping (3.10) or (3.11).
Lemma 3.2.
Suppose that (2.1) holds.
(3.12) $|m_{0}-1|\leq C|u_{0}|_{\mathfrak{M}^{0,0}},\ \ |\partial_{\lambda_{R}}m_{0}|,\ |\partial_{\lambda_{I}}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{0,0}}(1+|\lambda_{I}|),$
(3.13) $|\partial_{\lambda_{R}}^{j}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{j-1,0}}(1+|x_{2}|)^{j-1}(1+|\lambda_{I}|)^{j},\ j\leq 3,$
(3.14) $|\partial_{\lambda_{I}}^{l}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{1-l,0}}(1+|x_{2}|)^{l-1}(1+|\lambda|)^{l},\ 2\leq l\leq 3,$
(3.15) $|\partial_{\lambda_{R}}\partial_{\lambda_{I}}m_{0}|\leq C|u_{0}|_{\mathfrak{M}^{1,0}}(1+|x_{2}|)(1+|{\color[rgb]{0,0,0}\lambda}|)^{2}.$
Proof.
Estimate (3.12) follows from Lemma 3.1 and
(3.16) $\begin{gathered}m_{0}=(1-G_{u_{0}})^{-1}1,\\ \partial_{\lambda_{R}}m_{0}=(1-G_{u_{0}})^{-1}\left[\partial_{\lambda_{R}}G_{u_{0}}\right]m_{0},\quad\partial_{\lambda_{I}}m_{0}=(1-G_{u_{0}})^{-1}\left[\partial_{\lambda_{I}}G_{u_{0}}\right]m_{0}.\end{gathered}$
Applying the bootstrapping argument to (3.16) and Lemma 3.1, we can justify (3.13)-(3.15). ∎
Combining Lemma 3.2 with (2.2), (2.10), and (2.12), we further deduce
Lemma 3.3.
Suppose the assumption of Theorem 1 holds for $\mathfrak{M}^{2,q}$.
(3.17) $|\partial_{\lambda^{\prime}_{R}}^{j}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{j-1,0}}(1+|\lambda^{\prime}_{I}|)^{j},\ 1\leq j\leq 3,$
(3.18) $|\partial_{\lambda^{\prime}_{I}}^{l}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{l-1,0}}(1+|\lambda^{\prime}|)^{l},\ \ 1\leq l\leq 3,$
(3.19) $|\partial_{\lambda^{\prime}_{R}}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq$ $C|u_{0}|_{\mathfrak{M}^{1,0}}(1+|\lambda^{\prime}|)^{2}.$
@@ 3.2. Long time asymptotics of $u_{1}(x)$
Integration by parts, applying Lemmas 3.3, 3.4, we have
@@
Suppose the assumption of Theorem 1 holds for $\mathfrak{M}^{2,q}$. Then, as $t\to+\infty$:
@@
$=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\zeta^{\prime}_{I}e^{-2it(a\zeta^{\prime}_{I}-{\zeta^{\prime}_{I}}^{3})}e^{+\pi i\frac{\operatorname{sgn}(\zeta^{\prime}_{I})}{4}}\frac{1}{\sqrt{|\frac{6}{\pi}\zeta^{\prime}_{I}|}}\psi_{r,r}(\zeta^{\prime}_{I})(\overline{\zeta}^{\prime}-\zeta^{\prime})\widetilde{s}_{c}(0,\zeta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}}).$
@@
(3.49) $u_{1,1}(x)=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\eta^{\prime}_{I}\left(e^{-2it(a\zeta^{\prime}_{I}-{\zeta^{\prime}_{I}}^{3})}\right)^{\wedge_{\zeta^{\prime}_{I}}}(-\eta^{\prime}_{I})g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}})$
@@
Let $g_{+}=\psi_{r,0}(\zeta^{\prime}_{I})\partial_{\zeta^{\prime}_{R}}\left(\frac{1}{\zeta_{R}^{\prime}}\psi_{r,r}(\zeta^{\prime}_{R})\widetilde{s}_{c}(\zeta^{\prime})\right)$. Applying Lemma 3.3, and $u_{0}\in\mathfrak{M}^{1,q}$, we have
@@
Suppose that (2.1) holds for $\mathfrak{M}^{2,q}$. Then, as $t\to+\infty$,

### v4

@@ 3.1. Estimates on scattering data
(3.2) $\begin{gathered}G_{u_{0}}f=\iint e^{2\pi i(x_{1}\xi_{1}+x_{2}\xi_{2})}\frac{\left[u_{0}f\right]^{\wedge}(\xi_{1},\xi_{2},\lambda)}{p_{\lambda}(\xi_{1},\xi_{2})}d\xi_{1}d\xi_{2},\\ p_{\lambda}(\xi_{1},\xi_{2})=(2\pi i\xi_{1}+\lambda)^{2}-(2\pi i\xi_{2}+{\lambda}^{2}).\end{gathered}$
Applying the estimates [8]:
@@
where $\Omega_{\lambda}=\{(\xi_{1},\xi_{2})\in{\mathbb{R}}^{2}:|p_{\lambda}(\xi_{1},\xi_{2})|<1\}$, we have
(3.4) $|G_{u_{0}}f|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,0}}|f|_{L^{\infty}}.$
Moreover, via the Fourier theory and the residue theorem, the operator $G_{u_{0}}$ can be written as [1]:
@@
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left[4\pi i\xi_{1}(x_{2}-x_{2}^{\prime})\right]e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
${\color[rgb]{0,0,0}+}$ $\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi\lambda_{I})G_{u_{0}}(x_{2}-x_{2}^{\prime})f.$
As a result,
(3.7) $\begin{gathered}\partial_{\lambda_{R}}m_{0}=\sum_{k=0}^{1}\sum_{k^{\prime}=0}^{k}\lambda_{I}^{k}x_{2}^{k^{\prime}}m_{1,k,k^{\prime}},\quad|m_{1,k,k^{\prime}}|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{k-k^{\prime},0}},\end{gathered}$
and, for $\lambda^{\prime}_{I}\neq 0$,
(3.8) $|\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|=|\partial_{\lambda_{R}}s_{c}|$
$\leq$ $C|\lambda_{I}\left[x_{2}{u_{0}m_{0}}\right]^{\wedge}|+C|\left[{u_{0}m_{1,0,0}}\right]^{\wedge}|+C|\lambda_{I}\left[{u_{0}m_{1,1,0}}\right]^{\wedge}|+C|\lambda_{I}\left[{x_{2}u_{0}m_{1,1,1}}\right]^{\wedge}|$
$\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{1,0}})(1+|\lambda^{\prime}_{I}|).$
Lemma 3.1.
Suppose $|t_{2}|<C$ and $\lambda^{\prime}_{I}\neq 0$.
(3.9) $|\partial_{\lambda^{\prime}_{R}}^{j}\widetilde{s}_{c}|\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{j,0}})(1+|\lambda^{\prime}_{I}|)^{j},$
(3.10) $|\partial_{\lambda^{\prime}_{I}}^{j}\widetilde{s}_{c}|\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{j,0}})(1+|\lambda^{\prime}|)^{j},$
(3.11) $|\partial_{\lambda^{\prime}_{R}}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{2,0}})(1+|\lambda^{\prime}|)^{2}.$
Proof.
Similarly to the argument for (3.8), for $\lambda^{\prime}_{I}\neq 0$, using
(3.12) $\left[\partial_{\lambda_{R}}^{2}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
$\times$ $\theta((x_{2}-x_{2}^{\prime})\xi_{1}({\xi_{1}}+\frac{\lambda_{I}}{\pi}))\left(\frac{4\pi i}{-8\pi^{2}}\right)^{2}(4\pi\lambda_{I})(x_{2}-x_{2}^{\prime})\partial_{\xi_{1}}e^{-4\pi^{2}\xi_{1}(x_{2}-x_{2}^{\prime})({\xi_{1}}+\frac{\lambda_{I}}{\pi})}$
${\color[rgb]{0,0,0}+}$ $\left(\frac{4\pi i}{-8\pi^{2}}\right)(4\pi i\lambda_{I})\partial_{\lambda_{R}}\left(G_{u_{0}}(x_{2}-x_{2}^{\prime})f\right),$
$\vdots$
(3.13) $\left[\partial_{\lambda_{I}}G_{u_{0}}\right]f(x_{1},x_{2},\lambda)$
@@
$\vdots$
we have
(3.14) $\partial_{\lambda_{R}}\partial_{\lambda_{I}}m_{0}=$ $\sum_{k=0}^{2}\sum_{k^{\prime}=0}^{k}\lambda_{I}^{k}x_{2}^{k^{\prime}}m_{k,k^{\prime}},\quad|m_{k,k^{\prime}}|\leq C(1+|u_{0}|_{\mathfrak{M}^{k-k^{\prime},0}}),$
(3.15) $\partial^{j}_{\lambda_{R}}m_{0}=$ $\sum_{k=0}^{j}\sum_{k^{\prime}=0}^{k}\lambda_{I}^{k}x_{2}^{k^{\prime}}m_{j,k,k^{\prime}},\quad|m_{j,k,k^{\prime}}|\leq C(1+|u_{0}|_{\mathfrak{M}^{k-k^{\prime},0}}),$
(3.16) $\partial^{j}_{\lambda_{I}}m_{0}=$ ${\sum_{k=0}^{j}\sum_{k^{\prime}=0}^{k}\lambda_{I}^{k}x_{2}^{k^{\prime}}m^{+}_{j,k,k^{\prime}}}+\sum_{l=1}^{j-1}\sum_{h=0}^{j-l}\sum_{h^{\prime}=0}^{h}(x_{1}+\lambda_{R}x_{2})^{l}\lambda_{I}^{h}x^{h^{\prime}}_{2}m^{-}_{j,l,h,h^{\prime}},$
${|m^{+}_{j,k,k^{\prime}}|\leq C(1+|u_{0}|_{\mathfrak{M}^{k-k^{\prime},0}}),\quad|m^{-}_{j,l,h,h^{\prime}}|\leq C(1+|u_{0}|_{\mathfrak{M}^{l+h-h^{\prime},0}}).}$
Hence the proof of the lemma can be justified by taking derivatives of (2.2).
We have sharper estimates for the following first derivatives:
Lemma 3.2.
Suppose the assumption of Theorem 1 holds. For $\lambda^{\prime}_{I}\neq 0$,
(3.17) $|\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|\leq C{\color[rgb]{0,0,0}(1+|u_{0}|^{2}_{\mathfrak{M}^{1,1}}}),\quad|\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq C(1+\min\{|\lambda_{R}|,\frac{1}{|\lambda_{I}|}\}){\color[rgb]{0,0,0}(1+|u_{0}|^{2}_{\mathfrak{M}^{1,1}}}),$
(3.18) $|\lambda^{\prime}_{I}\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|\leq C{\color[rgb]{0,0,0}(1+|u_{0}|^{2}_{\mathfrak{M}^{1,2}}}),\quad|\lambda^{\prime}_{I}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq C(1+\min\{|\lambda_{R}|,\frac{1}{|\lambda_{I}|}\}){\color[rgb]{0,0,0}(1+|u_{0}|^{2}_{\mathfrak{M}^{1,2}}}).$
Proof.
For $j=1,2$, via (2.4), (3.1), the Fourier theory, and integration by parts,
(3.19) $|\left[\partial_{x_{j}}G_{u_{0}}\right]f|_{L^{\infty}}\leq C(|u_{0}|_{\mathfrak{M}^{0,1}}|f|_{L^{\infty}}+|u_{0}|_{\mathfrak{M}^{0,0}}|\partial_{x_{j}}f|_{L^{\infty}}),$
(3.20) $|\partial_{x_{j}}m_{0}|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,1}},$
(3.21) $|\left[\partial_{x_{j}}G_{u_{0}}\right]m_{0}|_{L^{\infty}}\leq C|u_{0}|_{\mathfrak{M}^{0,1}},$
(3.22) $|\partial_{x_{j}}m_{1,1,k}|_{L^{\infty}}\leq C(1+|u_{0}|_{\mathfrak{M}^{1,1}}),\quad k=0,1.$
Combining with (3.6) and integration by parts, for $\lambda^{\prime}_{I}\neq 0$, we obtain
(3.23) $|\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|=|\partial_{\lambda_{R}}s_{c}|$
$\leq$ $C(|\lambda_{I}\left[x_{2}{u_{0}m_{0}}\right]^{\wedge}|+|\left[{u_{0}m_{1,0,0}}\right]^{\wedge}|+|\lambda_{I}\left[{u_{0}m_{1,1,0}}\right]^{\wedge}|+|\lambda_{I}\left[{x_{2}u_{0}m_{1,1,1}}\right]^{\wedge}|)$
$\leq$ $C(|\left[\partial_{x_{1}}\{x_{2}{u_{0}m_{0}}\}\right]^{\wedge}|+|u_{0}|_{\mathfrak{M}^{0,0}}+|\left[\partial_{x_{1}}\{{u_{0}m_{1,1,0}}\}\right]^{\wedge}|+|\left[\partial_{x_{1}}\{{x_{2}u_{0}m_{1,1,1}}\}\right]^{\wedge}|)$
$\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{1,1}}),$
and, similarly,
(3.24) $|\lambda^{\prime}_{I}\partial_{\lambda^{\prime}_{R}}\widetilde{s}_{c}|\leq C|\lambda_{I}\partial_{\lambda_{R}}s_{c}|$
$\leq$ $C|\partial_{\lambda_{R}}\left[(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|+C|\partial_{\lambda_{R}}\left[u_{0}(\partial_{x_{1}}m_{0})\right]^{\wedge}|$
$\leq$ $C|\partial_{\lambda_{R}}\left[(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|+C|\partial_{\lambda_{R}}\left[u_{0}(1-G_{u_{0}})^{-1}(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|$
$\leq$ $C(1+|u_{0}|^{2}_{\mathfrak{M}^{1,2}}).$
In an entirely similar way, we can justify the corresponding estimates for $\partial_{\lambda_{I}^{\prime}}\widetilde{s}_{c}$ and $\lambda_{I}^{\prime}\partial_{\lambda_{I}^{\prime}}\widetilde{s}_{c}$:
(3.25) $|\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|=|\partial_{\lambda_{I}}s_{c}|$
$\leq$ $C|\left[x_{1}{u_{0}m_{0}}\right]^{\wedge}|+C|\lambda_{R}\left[x_{2}{u_{0}m_{0}}\right]^{\wedge}|$
[…]
$\leq$ $C|[x_{1}{u_{0}m_{0}}]^{\wedge}|+C\min\{\ |\lambda_{R}|\times|\left[x_{2}{u_{0}m_{0}}\right]^{\wedge}|,\ \frac{1}{|\lambda_{I}|}\times|\left[\partial_{x_{2}}\{x_{2}{u_{0}m_{0}}\}\right]^{\wedge}|\ \}$
[…]
$\leq$ $C(1+\min\{|\lambda_{R}|,\frac{1}{|\lambda_{I}|}\})(1+|u_{0}|^{2}_{\mathfrak{M}^{1,1}}),$
@@
(3.26) $|\lambda^{\prime}_{I}\partial_{\lambda^{\prime}_{I}}\widetilde{s}_{c}|\leq C|\lambda_{I}\partial_{\lambda_{I}}s_{c}|$
$\leq$ $C|\partial_{\lambda_{I}}\left[(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|+C|\partial_{\lambda_{I}}\left[u_{0}(\partial_{x_{1}}m_{0})\right]^{\wedge}|$
$\leq$ $C|\partial_{\lambda_{I}}\left[(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|+C|\partial_{\lambda_{I}}\left[u_{0}(1-G_{u_{0}})^{-1}(\partial_{x_{1}}u_{0})m_{0}\right]^{\wedge}|$
$\leq$ $C(1+\min\{|\lambda_{R}|,\frac{1}{|\lambda_{I}|}\})(1+|u_{0}|^{2}_{\mathfrak{M}^{1,2}}).$
@@ 3.2. Long time asymptotics of $u_{1}(x)$
Integration by parts, applying Lemmas 3.2, 3.3, we have
@@
Suppose the assumption of Theorem 1 holds. Then, as $t\to+\infty$:
@@
$=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\zeta^{\prime}_{I}e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3})}e^{+\pi i\frac{\operatorname{sgn}(\zeta^{\prime}_{I})}{4}}\frac{1}{\sqrt{|\frac{6}{\pi}\zeta^{\prime}_{I}|}}\psi_{r,r}(\zeta^{\prime}_{I})(\overline{\zeta}^{\prime}-\zeta^{\prime})\widetilde{s}_{c}(0,\zeta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}}).$
@@
(3.56) $u_{1,1}(x)=$ $-\frac{2}{\pi}\frac{1}{\sqrt{t}}\int d\eta^{\prime}_{I}\left(e^{-2it(a\zeta^{\prime}_{I}+{\zeta^{\prime}_{I}}^{3})}\right)^{\wedge_{\zeta^{\prime}_{I}}}(-\eta^{\prime}_{I})g^{\wedge_{\zeta^{\prime}_{I}}}(0,\eta^{\prime}_{I})+\mathcal{O}(\frac{1}{t^{\frac{3}{2}}})$
@@
Let $g_{+}=\psi_{r,0}(\zeta^{\prime}_{I})\partial_{\zeta^{\prime}_{R}}\left(\frac{1}{\zeta_{R}^{\prime}}\psi_{r,r}(\zeta^{\prime}_{R})\widetilde{s}_{c}(\zeta^{\prime})\right)$. Applying Lemma 3.1, and $\color[rgb]{0,0,0}u_{0}\in\mathfrak{M}^{2,q}$, we have
@@
Suppose that (2.1) holds. Then, as $t\to+\infty$,